OpenStudy (anonymous):

Help please!? Ms. Allen wants to give each of the 14 students in her class an equal number of pencils. The equation below can be used to find x, the total number of pencils she has to give her students. x/14 = 3 How many pencils does Ms. Allen have to give to her students? A. 11 B. 17 C. 32 D. 42

OpenStudy (anonymous):

x/14=3 Multiply both sides by 14. You get (14)(x/14)=(3)(14) which becomes (x)(14/14)=(3)(14) The fourteens divide out on the left hand side of the equation. So you now have x=(3)(14) Which reduces to x=42